OpenStudy (anonymous):

Solve: 6(x - 2) + 3(x - 2) = 18. @-------College Al…

OpenStudy (anonymous):

6x-12 + 3x-6 =18 9x-18=18 no solution i think lol

OpenStudy (anonymous):

wait no lol its x=4

OpenStudy (anonymous):

\[6(x-2)+3(x-2)=9(x-2)\] os you have \[9(x-2)=18\] \[x-2=2\] \[x=4\]

OpenStudy (anonymous):

x=4